OpenStudy (anonymous):

Help please!!....For what value of "k" does the equation kx^2 - 2kx + k - 1=0 not have solutions in the real number set?

mathslover (mathslover):

\[{p(x)=kx^2-2kx+k-1}\] let !

OpenStudy (anonymous):

for the value where k in in closed form

mathslover (mathslover):

factors of -1 = \(\pm 1\)

mathslover (mathslover):

p(-1)=\(\large{k(-1)^2-2k(-1)+k-1=0}\) \(\large{4k-1=0}\) \(\large{k=\frac{1}{4}}\)

mathslover (mathslover):

i am just confused ...

OpenStudy (across):

As long as \(k\geqslant0\), the polynomial will have real solutions.

jimthompson5910 (jim_thompson5910):

Hint: \[\Large ax^2+bx+c=0\] will have nonreal solutions when \[\Large b^2 - 4ac < 0\]

OpenStudy (anonymous):

make sure \(b^2-4ac\geq 0\) in this case \(a=k, b=-2k, c=(k-1)\)

OpenStudy (anonymous):

thanks for all the help guys!